# Robust iterative methods for linear systems with saddle point structure

### v1

@@ 8 $\ddagger$ $\ddagger$ 1 $3^{\star}$ $3^{\star}$ $7^{\star}$dagger $\dagger$
# $\tau_{large}$ $\tau_{mix}$ $\tau_{small}$ $10^{-5}$ $10^{-4}$ $10^{-3}$ $10^{-2}$ $10^{-1}$
1 21,512 1,065 37,526 $\dagger$ $\dagger$ $\dagger$ $\dagger$ $\dagger$
2 81,544 1,073 572,767 1,330,241 $\dagger$ $\dagger$ $\dagger$ $\dagger$
3 93,634 1,353 686,656 1,549,021 $\dagger$ $\dagger$ $\dagger$ $\dagger$
4 23,660 15,623 25,581 70,770 64,783 49,684 $\dagger$ $\dagger$
5 480,578 61,137 631,284 1,069,068 750,882 330,140 $\dagger$ $\dagger$
6 $92^{\ddagger}$ $44^{\ddagger}$ 93 119,534 96,486 67,828 $\dagger$ $\dagger$
7 2,698 1,502 4,203 22,391 17,769 14,439 8,017 $\dagger$
8 2,987 1,318$\ddagger$ 9,217 72,572 57,502 39,290 $\dagger$ $\dagger$
@@ 6 $92^{\ddagger}$ $44^{\ddagger}$ 93 119,534 96,486 67,828 $\dagger$ $\dagger$
Next, we consider the additional $M$-orthogonalization of the nullspace basis via the modified Gram-Schmidt procedure. The number of outer-inner iterations and the total number of nonzeros for the new scheme are presented in Tables 8, 9 and 10, respectively. In this case, the total number of nonzeros refers to the sum of the number of nonzeros of the nullspace basis, the $M$-orthogonalized nullspace basis and the nonzeros of the SPAI preconditioner (i.e., $nnz(\widetilde{Z})+nnz(\bar{Z})+nnz(\tilde{W})$). Although $\widetilde{Z}$ can be discarded after $\bar{Z}$ is computed, we include it in the sum as a measure of the peak memory requirement. When $\tau_{small}$ is used, there is not much difference in terms of inner and outer iterations, except for reorientation_8, where the number of average CG iterations drops from $20.5$ to $2.5$ if the nullspace basis is $M$-orthogonalized. When $\tau_{mix}$ is used, there are two problems for which the new scheme did not converge without $M$-orthogonalization, namely goddardRocketProblem_2 and tumorAntiAngiogenesis_8, these converge now with $M$-orthogonalization. GoddardRocketProblem_2, is particularly interesting, since it has a nullspace dimension of $1$, and hence $M$-orthogonalization is just normalizing the basis vector. The matrix $A$ has a condition number of $1.85\times 10^{15}$, i.e. it is almost singular, even though the condition number of the coefficient matrix $\mathcal{W}$ is $1.38\times 10^{7}$. In this case, the nullspace method becomes a more effective preconditioner with $M$-orthogonalization and the number of outer fGMRES iterations is reduced to $58$, without any increase in memory requirement. When $\tau_{large}$ is used, $M$-orthogonalization appears to cause more failures, because the nullspace basis is likely to be inaccurate to begin with, and further approximations only increase the error.
@@ 6.2 Generalized case
# $\tau_{large}$ $\tau_{mix}$ $\tau_{small}$ $10^{-5}$ $10^{-4}$ $10^{-3}$ $10^{-2}$ $10^{-1}$
1 3 2 2 3 $\dagger$ $\dagger$ $\dagger$ $\dagger$
2 2 2 2 2 $\dagger$ $\dagger$ $\dagger$ $\dagger$
3 2 1 1 2 4 6 12 13
4 2 2 2 $1^{\star}$ $2^{\star}$ 3 6 $32^{\star}$
5 2 3 2 $1^{\star}$ $1^{\star}$ 3 7 $44^{\star}$
6 3 3 1 $\dagger$ 2 3 7 $\ddagger$
7 3 3 2 $\dagger$ $\dagger$ $\ddagger$ $9^{\star}$ddagger
8 4 3 2 $\dagger$ $\dagger$ $\ddagger$ 10 $\dagger$
@@
In Table 12, we report the total number of nonzeros for the new scheme (i.e., $nnz(\widetilde{Z})+nnz(\widetilde{W})$) and ILUTP($\tau$) (i.e., $nnz(L)+nnz(U)$). The best method among those that converge without any failure for each problem is given in bold. The new scheme achieves fewer nonzeros compared to ILUTP for 3 cases, while ILUTP is better than the new scheme for 5 cases. However, in most cases, the difference in the number of nonzeros is not significant.
@@
# $\tau_{large}$ $\tau_{mix}$ $\tau_{small}$ $10^{-5}$ $10^{-4}$ $10^{-3}$ $10^{-2}$ $10^{-1}$
1 23,856 23,856 203,761 245,975 $\dagger$ $\dagger$ $\dagger$ $\dagger$
2 1,329 1,329 2,326 6,187 $\dagger$ $\dagger$ $\dagger$ $\dagger$
3 10,946 10,946 26,596 14,644 11,526 8,521 4,905 4,417
4 55,661 51,984 69,923 70,392 66,804 58,686 41,156 16,444
5 55,591 55,584 69,946 $72,739^{\star}$ 68,323 60,242 46,078 19,095
6 58,271 58,266 70,325 $\dagger$ 67,989 60,271 45,570 20,011
7 60,019 60,042 70,842 $\dagger$ $\dagger$ 64,229 46,681 $\dagger$
8 63,143 63,118 71,699 $\dagger$ $\dagger$ 62,842 45,280 $\dagger$
@@
Next, we consider the additional $M$-orthogonalization of the nullspace basis via the modified Gram-Schmidt procedure. The number of outer fGMRES(10) iterations and the total number of nonzeros for the new scheme are given in Tables 14 and 15, respectively. As before, the total number of nonzeros refers to the sum of the number of nonzeros of the nullspace basis, the $M$-orthogonalized nullspace basis and the nonzeros of the SPAI preconditioner (i.e. $nnz(\widetilde{Z})+nnz(\bar{Z})+nnz(\tilde{W})$). The average number of inner fGMRES(10), LSQR, and MRS are given in Table 16. For test problems in the generalized case, we do not observe any significant differences when $M$-orthogonalization is performed. There is some increase in the number of nonzeros in most cases.
@@ 6.3 General case
Finally, for the general problems given in Table 3, we present in Table 17 the number of outer GMRES(10) and fGMRES(10) iterations for the baseline method and the new scheme, respectively. The new scheme converges for all cases where $\tau_{mix}$ and $\tau_{small}$ are used, while ILUTP has at least one failure due to the encounter of a zero pivot for all drop tolerances. The new scheme fails in one case when $\tau_{large}$ is used due to reaching the maximum allowed iterations without reaching the required relative residual norm.
@@
# $\tau_{large}$ $\tau_{mix}$ $\tau_{small}$ $10^{-5}$ $10^{-4}$ $10^{-3}$ $10^{-2}$ $10^{-1}$
1 3 3 2 3 5 $\dagger$ $\dagger$ $\dagger$
2 10 9 2 4 9 $\dagger$ $\dagger$ $\dagger$
3 2 2 1 1 2 2 3 4
4 7 2 1 1 2 2 3 4
5 34 5 4 1 2 2 3 4
6 27 11 11 1 2 2 3 4
7 4 17 1 2 2 4 10 $\dagger$
8 $\ddagger$ 53 5 $\dagger$ $\dagger$ $\dagger$ $\dagger$ $\dagger$
9 9 2 2 1 1 $\dagger$ $\dagger$ $\dagger$
@@ 8 $\ddagger$ 53 5 $\dagger$ $\dagger$ $\dagger$ $\dagger$ $\dagger$
In Table 18, we report the total number of nonzeros for the new scheme (i.e., $nnz(\widetilde{Z})+nnz(\widetilde{U})+nnz(\widetilde{W})$) and ILUTP($\tau$) (i.e., $nnz(L)+nnz(U)$). The best method among those that converge without any failure for each problem is given in bold. The new scheme not only converges without any failures for all problems but also achieves fewer nonzeros compared to ILUTP for all cases, except for two (garon_1 and garon_2). We note that for these two cases, even though ILUTP achieves fewer number of nonzeros, it is likely to fail if the drop tolerance is not chosen correctly.
@@
# $\tau_{large}$ $\tau_{mix}$ $\tau_{small}$ $10^{-5}$ $10^{-4}$ $10^{-3}$ $10^{-2}$ $10^{-1}$
1 1,252,275 1,252,275 2,204,857 644,190 623,798 $\dagger$ $\dagger$ $\dagger$
2 18,037,895 18,037,895 31,745,441 5,502,974 5,345,750 $\dagger$ $\dagger$ $\dagger$
3 162 162 162 1,829 1,797 1,642 1,229 640
4 612 612 612 2,529 2,497 2,342 1,860 1,205
5 1,962 1,962 1,962 4,629 4,597 4,442 3,618 2,882
6 7,200 7,200 7,200 12,777 12,745 11,603 10,233 9,344
7 1,885 1,688 1,875 28,035 27,955 27,527 25,366 $\dagger$
8 114114$\ddagger$114 ‡ 90 134 $\dagger$ $\dagger$ $\dagger$ $\dagger$ $\dagger$
9 320 320 320 5,298 5,298 $\dagger$ $\dagger$ $\dagger$

### v2

@@ 8 $\ddagger$ $\ddagger$ 1 $3^{\star}$ $7^{\star}$ $\dagger$ $\dagger$ 3
# $\tau_{large}$ $\tau_{mix}$ $\tau_{small}$ $10^{-4}$ $10^{-3}$ $10^{-2}$ $10^{-1}$ sym-ildl
1 21,512 1,065 37,526 $\dagger$ $\dagger$ $\dagger$ $\dagger$ 3,639
2 81,544 1,073 572,767 $\dagger$ $\dagger$ $\dagger$ $\dagger$ 52,583
3 93,634 1,353 686,656 $\dagger$ $\dagger$ $\dagger$ $\dagger$ 58,407
4 23,660 15,623 25,581 64,783 49,684 $\dagger$ $\dagger$ 9,626
5 480,578 61,137 631,284 750,882 330,140 $\dagger$ $\dagger$ 55,834
6 $92^{\ddagger}$ $44^{\ddagger}$ 93 96,486 67,828 $\dagger$ $\dagger$ 14,535
7 2,698 1,502 4,203 17,769 14,439 8,017 $\dagger$ 3,131
[…]
@@ 6 $92^{\ddagger}$ $44^{\ddagger}$ 93 96,486 67,828 $\dagger$ $\dagger$ 14,535
Next, we consider the additional $M$-orthogonalization of the nullspace basis via the modified Gram-Schmidt procedure. The number of outer-inner iterations and the total number of nonzeros for the new scheme are presented in Tables 8, 9 and 10, respectively. In this case, the total number of nonzeros refers to the sum of the number of nonzeros of the nullspace basis, the $M$-orthogonalized nullspace basis and the nonzeros of the SPAI preconditioner (i.e., nnz(Z~)+nnz(Z¯)+nnz(W~)nnz($\widetilde$$\bar{Z})+nnz(\tilde${W})). Although $\widetilde{Z}$ can be discarded after $\bar{Z}$ is computed, we include it in the sum as a measure of the peak memory requirement. When $\tau_{small}$ is used, there is not much difference in terms of inner and outer iterations, except for reorientation_8, where the number of average CG iterations drops from $20.5$ to $2.5$ if the nullspace basis is $M$-orthogonalized. When $\tau_{mix}$ is used, there are two problems for which the new scheme did not converge without $M$-orthogonalization, namely goddardRocketProblem_2 and tumorAntiAngiogenesis_8, these converge now with $M$-orthogonalization. GoddardRocketProblem_2, is particularly interesting, since it has a nullspace dimension of $1$, and hence $M$-orthogonalization is just normalizing the basis vector. The matrix $A$ has a condition number of $1.85\times 10^{15}$, i.e. it is almost singular, even though the condition number of the coefficient matrix $\mathcal{W}$ is $1.38\times 10^{7}$. In this case, the nullspace method becomes a more effective preconditioner with $M$-orthogonalization and the number of outer fGMRES iterations is reduced to $58$, without any increase in memory requirement. When $\tau_{large}$ is used, $M$-orthogonalization appears to cause more failures, because the nullspace basis is likely to be inaccurate to begin with, and further approximations only increase the error.
@@ 6.2 Generalized case
# $\tau_{large}$ $\tau_{mix}$ $\tau_{small}$ $10^{-4}$ $10^{-3}$ $10^{-2}$ $10^{-1}$
1 3 2 2 $\dagger$ $\dagger$ $\dagger$ $\dagger$
2 2 2 2 $\dagger$ $\dagger$ $\dagger$ $\dagger$
3 2 1 1 4 6 12 13
4 2 2 2 $2^{\star}$ 3 6 $32^{\star}$
5 2 3 2 $1^{\star}$ 3 7 $44^{\star}$
6 3 3 1 2 3 7 $\ddagger$
7 3 3 2 $\dagger$ $\ddagger$ $9^{\star}$ $\ddagger$
8 4 3 2 $\dagger$ $\ddagger$ 10 $\dagger$
@@
In Table 12, we report the total number of nonzeros for the new scheme (i.e., nnz(Z~)+nnz(W~)nnz($\widetilde${Z})+nnz($\widetilde${W})) and ILUTP($\tau$) (i.e., $nnz(L)+nnz(U)$). The best method among those that converge without any failure for each problem is given in bold. The new scheme achieves fewer nonzeros compared to ILUTP for 3 cases, while ILUTP is better than the new scheme for 5 cases. However, in most cases, the difference in the number of nonzeros is not significant.
@@
# $\tau_{large}$ $\tau_{mix}$ $\tau_{small}$ $10^{-4}$ $10^{-3}$ $10^{-2}$ $10^{-1}$
1 23,856 23,856 203,761 $\dagger$ $\dagger$ $\dagger$ $\dagger$
2 1,329 1,329 2,326 $\dagger$ $\dagger$ $\dagger$ $\dagger$
3 10,946 10,946 26,596 11,526 8,521 4,905 4,417
4 55,661 51,984 69,923 66,804 58,686 41,156 16,444
5 55,591 55,584 69,946 68,323 60,242 46,078 19,095
6 58,271 58,266 70,325 67,989 60,271 45,570 20,011
7 60,019 60,042 70,842 $\dagger$ 64,229 46,681 $\dagger$
8 63,143 63,118 71,699 $\dagger$ 62,842 45,280 $\dagger$
@@
Next, we consider the additional $M$-orthogonalization of the nullspace basis via the modified Gram-Schmidt procedure. The number of outer fGMRES(10) iterations and the total number of nonzeros for the new scheme are given in Tables 14 and 15, respectively. As before, the total number of nonzeros refers to the sum of the number of nonzeros of the nullspace basis, the $M$-orthogonalized nullspace basis and the nonzeros of the SPAI preconditioner (i.e. nnz(Z~)+nnz(Z¯)+nnz(W~)nnz($\widetilde$$\bar{Z})+nnz(\tilde${W})). The average number of inner fGMRES(10), LSQR, and MRS are given in Table 16. For test problems in the generalized case, we do not observe any significant differences when $M$-orthogonalization is performed. There is some increase in the number of nonzeros in most cases.
@@ 6.3 General case
For the general problems given in Table 3, in Table 17 we present the number of outer GMRES(10) and fGMRES(10) iterations for the baseline method and the new scheme, respectively. The new scheme converges for all cases where $\tau_{mix}$ and $\tau_{small}$ are used, while ILUTP has at least one failure due to the encounter of a zero pivot for all drop tolerances. The new scheme fails in one case when $\tau_{large}$ is used due to reaching the maximum allowed iterations without reaching the required relative residual norm.
@@
# $\tau_{large}$ $\tau_{mix}$ $\tau_{small}$ $10^{-4}$ $10^{-3}$ $10^{-2}$ $10^{-1}$
1 3 3 2 5 $\dagger$ $\dagger$ $\dagger$
2 10 9 2 9 $\dagger$ $\dagger$ $\dagger$
3 2 2 1 2 2 3 4
4 7 2 1 2 2 3 4
5 34 5 4 2 2 3 4
6 27 11 11 2 2 3 4
7 4 17 1 2 4 10 $\dagger$
8 $\ddagger$ 53 5 $\dagger$ $\dagger$ $\dagger$ $\dagger$
9 9 2 2 1 $\dagger$ $\dagger$ $\dagger$
@@ 8 $\ddagger$ 53 5 $\dagger$ $\dagger$ $\dagger$ $\dagger$
In Table 18, we report the total number of nonzeros for the new scheme (i.e., nnz(Z~)+nnz(U~)+nnz(W~)nnz($\widetilde${Z})+nnz($\widetilde${U})+nnz($\widetilde${W})) and ILUTP($\tau$) (i.e., $nnz(L)+nnz(U)$). The best method among those that converge without any failure for each problem is given in bold. The new scheme not only converges without any failures for all problems but also achieves fewer nonzeros compared to ILUTP for all cases, except for two (garon_1 and garon_2). We note that for these two cases, even though ILUTP achieves fewer number of nonzeros, it is likely to fail if the drop tolerance is not chosen correctly.
@@
# $\tau_{large}$ $\tau_{mix}$ $\tau_{small}$ $10^{-4}$ $10^{-3}$ $10^{-2}$ $10^{-1}$
1 1,252,275 1,252,275 2,204,857 623,798 $\dagger$ $\dagger$ $\dagger$
2 18,037,895 18,037,895 31,745,441 5,345,750 $\dagger$ $\dagger$ $\dagger$
3 162 162 162 1,797 1,642 1,229 640
4 612 612 612 2,497 2,342 1,860 1,205
5 1,962 1,962 1,962 4,597 4,442 3,618 2,882
6 7,200 7,200 7,200 12,745 11,603 10,233 9,344
7 1,885 1,688 1,875 27,955 27,527 25,366 $\dagger$
8 $114\ddagger$ 90 134 $\dagger$ $\dagger$ $\dagger$ $\dagger$
9 320 320 320 5,298 $\dagger$ $\dagger$ $\dagger$